OpenStudy (anonymous):

evaluate the given integral using a substitution (xe^-x^2+8 +e^5x) dx

OpenStudy (anonymous):

\[\int xe^{-x^2}\;dx+\int 8\;dx+\int e^{5x}\;dx\] For the first integral, let \(u=-x^2\). For the third, let \(u=5x\). The second integral doesn't require a substitution.